OpenStudy (anonymous):

Ms. Carrs class is selling magazines to raise money for a field trip. the students in the class want to decide how much profit they'll make from each magazine they sell. Rewrite the formula to determine the number of magazinees they'll need to sell

OpenStudy (anonymous):

the following formula describes the sceanario: p= m/n p=profit ade from each magazine sold m=money they want to raise n=number of magazines sold

OpenStudy (anonymous):

m=n*p? maybe?

OpenStudy (anonymous):

wrong..... thanks anyways